### v2

@@ Abstract
Over the last five years, McGill University’s Office of Science Education (OSE) has partnered with faculty members from the Department of Physics to form an education working group with the aim of charting the progression of students’ conceptual understanding of uncertainties across their undergraduate degree. The research conducted by this group seeks to provide further insight into both the experimental skillset that students gain through undergraduate laboratory courses and how the department could address noticeable gaps in student understanding. In this paper, we evaluate the conceptual understanding of uncertainty using the Concise Data Processing Assessment (CDPA) test. First, we characterize the physics laboratory curriculum at McGill University by evaluating the evolution of CDPA scores across consecutive laboratory courses, and further propose the utilization of this tool for identifying of gaps in student understanding. Following the analysis of student responses (N = 2023) we specifically investigate data collected in second-year courses to better understand what student errors reveals about common misconceptions in experimental physics. This more in-depth analysis focuses on data collected from students at the beginning and end of two consecutive experimental laboratory courses that build on each other. By the end of the second course, students have engaged with all the material covered in the CDPA test. Interestingly, we did not observe any change in CDPA scores throughout and post pandemic. We notice a marked upward shift in student understanding of uncertainties, however the results indicate that many students continue to struggle with uncertainties, basic data analysis and curve fitting.
@@ I Introduction
Understanding measurements of physical quantities and uncertainties is an important skillset for evaluating and interpreting data in scientific research. In recent years, the consideration and discussion of uncertainties has increasingly become more present and relevant in physics undergraduate curricula [1, 2, 3], in line with the American Association of Physics Teachers (AAPT) recommendations for undergraduate physics labs [4]. Skills such as modeling [5], experimental design [6], data analysis, interpretation [7], and communication [8] are essential for the training of physicists. Developing a comprehensive understanding of how these competencies are scaffolded across an educational curricula can guide instructional changes and reinforce pedagogical practices that are further in line with the AAPT guidelines.
The study reported in this paper was conducted at McGill University, a research-intensive university in Quebec, Canada. Following an in-depth curriculum review process, the Chair of Physics Department invited a small group of professors teaching at different levels in the experimental physics program to partner with the University’s Office of Science Education (OSE) to form the physics education research (PER) group. Improving introductory, intermediate, and advanced undergraduate physics labs to reinforce the understanding of uncertainties was selected as a central point. This focus emerged through in-depth discussions about assessment conducted at peer institutions [9, 10, 11, 12, 13]. These studies highlight the importance of assessing the understanding of uncertainties as a means to revise and improve upon students’ experimental competencies in first-year laboratory courses [14, 15].
In the PER literature, a range of tests are applied to assess experimental competencies such as the Concise Data Processing Assessment (CDPA) [10], the Colorado Learning Attitudes about Science Survey for Experimental Physics (E-CLASS) [16], the Physics Measurement Questionnaire (PMQ) [17], the Laboratory Data Analysis Instrument (LDAI)[18], the Physics Lab Inventory of Critical Thinking (PLIC) [19], and more recently the Survey of Physics Reasoning on Uncertainty Concepts in Experiments (SPRUCE) [20]. Each of these evaluates different aspects of physics students’ learning. PLIC, for example, focuses on students’ evidence-based decision making by measuring abilities ranging from evaluation of data, methods, and conclusions to the proposal of next steps in hypothetical experiments. In comparison, SPRUCE and CDPA are examples of specific tools to assess the understanding of the concept of uncertainty in undergraduate labs. SPRUCE has a broader approach, identifying sources of uncertainty while an experiment is being conducted and leading up to the handling of uncertainty, whereas the CDPA is used specifically to assess error propagation skills, and the interpretation of equations, graphs, and uncertainties therein.
@@
The objectives of this work are 1) to establish a validation of baseline CDPA data at McGill University compared to the data published in a previous study [10], 2) to investigate the progression of CDPA scores throughout different lab courses and across consecutive years, 3) to compare the evolution of student comprehension on the topic of uncertainty prior to, during, and directly following the global COVID-19 pandemic period (2019-2022) during which students experienced varying degrees of virtual versus in-person instruction, and finally 4) to propose a deeper interpretation of CDPA results in terms of mapping student abilities and identifying common student misconceptions. Where possible, we also provide additional contextual explanations of the features of our findings that may influence the data interpretation.
@@ II Background
The CDPA was first implemented in 2011 at the University of British Columbia (UBC) and the University of Edinburgh [10]. It was initially developed for first-year physics major students in laboratory courses or in science-intensive programs that also required an experimental physics laboratory component. Graduate students and professors also participated in the study in order to ensure the validity of the test in measuring novice vs. expert performance and to further investigate its reliability when used by students at different levels of instruction. The CDPA test consists of 10 multiple choice questions, which vary in terms of difficulty, each with four possible answers. The questions cover a variety of learning outcomes such as identifying the adequacy of a fit to a given dataset, handling instrumental uncertainty, and inferring measurement units from the data. The test’s characteristics, such as its difficulty and its capacity to discriminate amongst the level of expertise of test-takers, are proposed and quantified in the original paper. Furthermore, the test is administered both as a pre-course and post-course assessment in order to reliably measure shifts in reasoning that can be linked to the learning outcomes of each respective course.
@@
The results in [10] showed 1) that there are positive shifts in pre- to post- comparisons for the first-year courses, 2) that the test is sensitive to the level of expertise of the test-taker, and thus is able to distinguish between undergraduate students, graduate students, and professors, 3) that the test is suited to a broader public, a conclusion made based on the comparable results from UBC and the University of Edinburgh. All these findings are supported by statistical tests which confirm the reliability of the CDPA. For example, first-year UBC students scored an average of $3.91\pm 0.12$ (standard error of the mean, SEM) in the post-test, compared to an average of $5.31\pm 0.42$ (SEM) for the graduate students, and $8.00\pm 0.35$ (SEM) for professors. A similar, more marginal progression of knowledge was also shown across the undergraduate program by comparing the post-test of first-year students with that of fourth-year students who averaged $4.73\pm 0.26$ (SEM) in the assessment.
@@
Since its publication, this ten-question test has stimulated discussions amongst the PER community with different proposed applications of the CDPA results. Namely, Day et al. [21] investigated the influence of gender in the assessment results, finding no evidence of gender differences in conceptual understanding of uncertainties. The CDPA was also the instrument used to quantify the role of guidance during the learning process and it was found that scaffolding student orientation improves performance compared to students that do not receive any guidance, and that there is a link between the invention process and performance [22]. Kontro, et al. [23] further explore the use of the CDPA for intermediate-level undergraduate physics courses and proposes the possibility of a relationship between CDPA final scores and E-CLASS results, though no significant correlation is observed.
@@ III.1 Context
The findings presented in this paper span the academic years from 2019 to 2023. In response to the global pandemic in 2020, universities were forced to adopt an remote learning. Following this period, 2021 marked the beginning of a hybrid instructional format with classes held online, while lab sessions gradually transitioned back to an in-person format. In 2022, the laboratory classes covered in this study returned to a fully in-person format. For the purposes of this study, we consider that the contents of these courses did not change between the in-person to the online setting, since the courses followed the same learning outcomes in both cases. In keeping with the CDPA requirements, no adaptations were made to the original CDPA test and the entire ten-question test was provided to all student respondents. Furthermore, as requested by the authors, we did not reproduce or include direct citations of any of the ten CDPA questions in order to preserve the possible use of this assessment instrument for future research purposes. However, it must also be noted that, in order to ensure a higher proportion of student respondents, a bonus of $1\%$ in their final course grade was offered in each class if at least $80\%$ of students in a given course took the test.
Finally, to avoid statistically significant shifts due to students retaking the CDPA test, we made sure to ask student respondents to identify whether or not they had taken the CDPA test in a previous semester. Students with prior exposure to the CDPA test were omitted from this study. The test was conducted online through MS Forms and student respondents were instructed to use up to 30 minutes to complete it. The pre-tests were taken over the first three to four weeks of term, whereas post-tests were taken during the last three to four weeks of term. Tests without pre and post distincion were taken at the end of the term. We investigated a total of 8 courses: PHYS 101- Introduction to Mechanics, PHYS 131- Mechanics and Waves, PHYS 257- Experimental Methods I, PHYS 258- Experimental Methods II, PHYS 359- Honors Laboratory in Modern Physics I, PHYS 439- Majors Laboratory in Modern Physics, PHYS 469- Honors Laboratory in Modern Physics 2, and PHYS 434- Optics. All results presented here are summarized in Appendix appendix.
@@ III.2 McGill Physics curriculum
The understanding of uncertainty in McGill Physics laboratory courses becomes progressively more important as the level of lab increases. This increasing importance is reflected in the progression of the weight assigned to students’ abilities to interpret and display their understanding of uncertainties in reports and graphs as they move closer to the end of their degree.
First-year students are often exposed to experimental physics through a laboratory portion of an introductory course if they are enrolled in a general science program or are pursuing a minor in physics. In these cases, at McGill University, students are required to take an introductory undergraduate laboratory course: PHYS 101- Introduction to Mechanics. The alternative for physics majors or honors students is to take part in the higher-level introductory course: PHYS 131- Mechanics and Waves. In these introductory courses, students explore basic mechanics concepts through simple confirmation experiments, where the concept of uncertainty is addressed but does not constitute a major portion of their final grade. Students who proceed to a physics-specific degree, i.e., majors and honors students, enroll during their second year in PHYS 257- Experimental Methods I and PHYS 258- Experimental Methods II, where they are formally introduced to uncertainties in experiments. In these two courses, which span one full academic year, students are required to report uncertainties and discuss the validity of their experimental results and findings based on their interpretation of these uncertainties.
By the conclusion of PHYS 257 and PHYS 258, students are expected to possess the skills needed to interpret, address, and visually represent measurements. Students who choose to focus on experimental physics take the 300- and 400-level courses, which are not part of the mandatory curriculum, and can therefore represent expertise-like level of understanding. The notation difference between 300 and 400 is simply a means to indicate which students are part of the honors program as opposed to the majors program, otherwise the curricular content covered in both PHYS 359 and PHYS 439 is identical. We did not consider a chart uniquely for PHYS 469, PHYS 434 and PHYS 439. We decided to join the PHYS 434, PHYS 439 and PHYS 469 courses to form the overall 400 level, because the sample size was too small and they correspond to similar levels of expertise.
@@ III.3 Educational background
In the context of our university, situated in Québec, a notable portion of incoming freshman students exhibit diverse educational backgrounds. This diversity is attributed to the fact that some students have completed Collège d’Enseignement Général et Professionnel (CEGEP) before enrolling in university. Students who complete CEGEP and pursue a physics degree at univeristy skip the first year of the McGill curriculum, taking PHYS 257 and PHYS 258 as their initial university-level laboratory courses. During data collection and analysis, we could not discriminate whether students had completed CEGEP.
@@ III.4.1 Origin of the PER group
The PER group who led for this study represented a partnership between the Department of Physics and the Office of Science Education (OSE). Following an in-depth curriculum review process, the chair of the department nominated five tenure track professors teaching at different levels in the experimental physics program to work with OSE staff (including the director, educational developers and a science education expert), the physics liaison librarian, as well as PhD students selected by the department. The goal was to measure the effectiveness of lab courses from a program perspective. Although membership evolved over the five-year period, the group met on a regular basis and operated similarly to a community of practice [cuddy2002cultivating] with distributed leadership and regular cycles of data collection and reflection.
@@ III.4.2 PER professors feedback sessions
Following the most recent round of data collection, feedback sessions were held with the PER professors. The purpose of these sessions was to better understand the level of each of the department’s experimental courses and to map the specific abilities taught in each course as they related to the content of the CDPA questions. These sessions were used to situate each course in the McGill Physics undergraduate program based on their contents, and to better understand how students at various points in their degrees would be expected to respond to the CDPA. This analysis will be discussed further below.
@@ IV.1 Comparison to UBC
Establishing a comparison between universities across the country is useful to evaluate the progression of students at each level of their studies across academic years. In Figure 1, we present the comparisons between the results at McGill University (in blue) and the University of British Columbia (in blue), where the CDPA survey was first developed. At McGill, we consider the results from the year of 2019, comparing it to the reported results at UBC. The overall results across both universities showed an ability differentiate between novice and expert-level understanding as a function of a respondent’s level of education. We notice a maximum score for students in the fourth and final year of their undergraduate education, without significant change towards the graduate student level. As expected, professors from both universities score higher than all previous cohorts, but not a perfect score, which also points to to the high standard of difficulty of this test. These CDPA results allow us to draw an important conclusion: even at the higher end of the novice to expert scale, the average score is not a complete $10$, a feature consistent with Day et. al. [10].
@@ IV.2 Longitudinal analysis of the McGill data
The total CDPA mean score for the courses investigated between 2019-2023 is presented on a scale from 0 to 10 in Figure 2a. From left to right, the order of level of expertise increases by course numbers. We notice statistically significant jumps in the CDPA scores when comparing the scores obtained during the very first mechanics course that contains a lab component, PHYS 101, to a more advanced laboratory, PHYS 257, when students are properly exposed to the concepts of uncertainty and the methods used to analyze them experimentally.
@@
In Figure 2b, we compare how the CDPA scores evolve for the same course as a function of time. Interestingly, for PHYS 101 and PHYS 131, the overall CDPA score has a combined average that does not change as a function of year, except in the rare case of PHYS 101 in 2022 where the score was abnormally low. It is important to note that no major changes were implemented in the course that could have explained this difference in score. In PHYS 257 and PHYS 258, the results show fluctuations that can be consistent with the learning curve faced at that level for each of the different cohorts. The structure, contents and evaluation methods used in these two courses did not change in any of the years under study. However, the difference is that in each year different instructors taught the course using the same syllabus . In the 2020 Winter session of PHYS 258, students only completed the laboratory component of the class which corresponds to 60 % of the syllabus, not accounting for the final project. That may partially explain the decrease in CDPA scores that year.
For higher level courses over the first three years, the overall score is also consistent with a jump observed in 2022. In that year, the number of students taking part in the class was smaller than in previous years but students still obtained significantly higher scores than previous cohorts. These findings are likely due to classroom dynamics that are notably different from the 200 level courses, which we will further argue plays an important role in students’ understanding of uncertainty.
@@
The data presented in Figure 2 shows that for the two lower-level courses, PHYS 101 and PHYS 131, the overall CDPA score can be understood as being comparable to random guessing. This can be explained by the fact that, in those courses, students are introduced to the concept of uncertainties for the first time and are only capable of applying simple mechanics problems and experiments without a deeper conceptual understanding. The CDPA test evaluates student understanding in topics that involve uncertainty propagation, graph analysis with fits and fitting uncertainties, logarithmic fits, and evaluation of $\chi^{2}$. In the original CDPA paper [10], the authors discuss that the vast majority of students from introductory-level laboratory courses are not yet ready to answer all the survey questions in the first year of their undergraduate degree. This portrayal of the perceived accuracy of the CDPA measurement tool for introductory-level courses is in agreement with our observations and analysis of the understanding of uncertainty for students enrolled in PHYS 101 and PHYS 131.
In the following discussions, we hone in on the results obtained from the PHYS 257 and PHYS 258 courses because it is during this full academic year that students learn how to perform curve fitting, start considering uncertainties to assess the outcome of an experiment, and produce and interpret graphs based on measuring uncertainties. Comparisons at the higher level could also garner interesting results, however the population size of student respondents was not sufficiently large for each year of the study. Additionally, having pairwise comparisons between two courses within the same academic year allows for a better comparison in terms of the understanding of uncertainties. In the case of the McGill Physics curriculum, the two courses are complementary to one another and students’ abilities to understand concepts related to uncertainty is not developed until the end of the full academic year, as shown in Table 1. At the end of PHYS 258, students are asked to conduct a small research project in which they need to answer a research question of their choice. Students experimentally validate their measurements and study how the uncertainties should be considered in the analysis, as well as how to best mitigate errors.
@@
IV.3 Pandemic results
The global COVID-19 pandemic of 2020 introduced a significant external factor that is separately considered in this work. As was earlier stated, although the syllabi of the courses included in this study did not change before, during, and after the pandemic, the formats in which these courses were delivered had to adapt due to provincial regulatory health guidelines. As such, even after smaller courses returned to an in-person teaching environment, many of the large-scale introductory courses, including PHYS 101, PHYS 131, PHYS 257, and PHYS 258 continued to be taught in either a hybrid or a fully remote format.
Surprisingly, however, we notice no significant difference in CDPA scores between these pre- and post-pandemic periods. As was pointed out by one of the course instructors, this lack of any statistically significant fluctuation in CDPA scores is, in and of itself, noteworthy. The consistency of these scores, as shown in Figures 2 and Figure 3, suggests that the pedagogical benefit of laboratory courses may have more to do with students gaining first-hand practical knowledge through manipulating laboratory instruments than with their conceptual understanding of measurement principles.
Furthermore, if exposure to experimental instruments is only part of the intended pedagogical value of laboratory courses, then how can these in-person labs be improved such that they enhance students’ understanding of concepts related to uncertainty to a degree that is distinguishable using a measurement instrument such as the CDPA? The consistency of these pandemic-era CDPA scores also raises the question whether the CDPA instrument can to distinguish between the impact of confirmation versus inquiry-guided instructional approaches. However, from the findings of this research, we deduce that in all undergraduate courses included in this study could benefit from more dedicated teaching of graphing and analysis of uncertainties, as we will discuss in the following sections.

### v3

@@ Abstract
Over the last five years, McGill University’s Office of Science Education (OSE) has partnered with faculty members from the Department of Physics to form an education research group with the aim of charting the progression of students’ conceptual understanding of uncertainties across their undergraduate degree. The research conducted by this group seeks to provide further insight into the experimental skillset that students gain through undergraduate laboratory courses and noticeable gaps in student understanding that the department could address. In this paper, we evaluate the conceptual understanding of uncertainty using the Concise Data Processing Assessment (CDPA) test. First, we characterize the physics laboratory curriculum at McGill University by evaluating the evolution of CDPA scores across consecutive laboratory courses, and further propose the utilization of this tool for identifying gaps in student understanding. Following the analysis of student responses (N = 2023) we specifically investigate data collected in second-year courses to better understand what student errors reveal about common misconceptions in experimental physics. This more in-depth analysis focuses on data collected from students at the beginning and end of two consecutive experimental laboratory courses that build on each other. By the end of the second course, students have engaged with all the material covered in the CDPA test. Overall results show an upward shift in student understanding of uncertainties over time. Interestingly, we did not observe any change in CDPA scores comparing throughout and post pandemic scores. Despite the upward shift, many students continue to struggle with uncertainties, basic data analysis and curve fitting.
@@ I Introduction
Understanding measurements of physical quantities and uncertainties is an important skillset for evaluating and interpreting data in scientific research. In recent years, the consideration and discussion of uncertainties has increasingly become more present and relevant in physics undergraduate curricula [1, 2, 3], in line with the American Association of Physics Teachers (AAPT) recommendations for undergraduate physics labs [4]. Skills such as modeling [5], experimental design [6], data analysis, interpretation [7], and communication [8] are essential for the training of physicists. Developing an understanding of where the competencies are not being addressed within the physics curriculum can help professors develop pedagogical practices to address these gaps.
The study reported in this paper was conducted at McGill University, a research-intensive university in Quebec, Canada. Following an in-depth curriculum review process, the Chair of Physics Department invited a small group of professors teaching at different levels in the experimental physics program to partner with the University’s Office of Science Education (OSE) to form the physics education research (PER) group. The focus of the research was chosen to be improving introductory, intermediate, and advanced undergraduate physics labs to reinforce the understanding of uncertainties was selected the focus of our research. This focus emerged through in-depth discussions about assessment conducted at peer institutions [9, 10, 11, 12, 13]. These studies highlight the importance of assessing the understanding of uncertainties as a means to revise and improve students’ experimental competencies in first-year laboratory courses [14, 15].
In the PER literature, a range of tests are used to assess experimental competencies such as the Concise Data Processing Assessment (CDPA) [10], the Colorado Learning Attitudes about Science Survey for Experimental Physics (E-CLASS) [16], the Physics Measurement Questionnaire (PMQ) [17], the Laboratory Data Analysis Instrument (LDAI)[18], the Physics Lab Inventory of Critical Thinking (PLIC) [19], and more recently the Survey of Physics Reasoning on Uncertainty Concepts in Experiments (SPRUCE) [20]. Each of these evaluates different aspects of physics students’ learning. PLIC, for example, focuses on students’ evidence-based decision making by measuring abilities ranging from evaluation of data, methods, and conclusions to the proposal of next steps in hypothetical experiments. In comparison, SPRUCE and CDPA are examples of specific tools to assess the understanding of the concept of uncertainty in undergraduate labs. SPRUCE has a broader approach, identifying sources of uncertainty while an experiment is being conducted and leading up to the handling of uncertainty, whereas the CDPA is used specifically to assess error propagation skills, and the interpretation of equations, graphs, and uncertainties therein.
@@
The objectives of this work are 1) to establish a validation of baseline CDPA data at McGill University compared to the data published in a previous study [10], 2) to investigate the progression of CDPA scores throughout different lab courses in consecutive years, 3) to compare the evolution of student comprehension on the topic of uncertainty prior to, during, and directly following the global COVID-19 pandemic period (2019-2022) during which students experienced varying degrees of virtual versus in-person instruction, and finally 4) to propose a deeper interpretation of CDPA results in terms of mapping student abilities and identifying common student misconceptions. Where possible, we also provide additional contextual explanations of the features of our findings that may influence the data interpretation.
@@ II Background
The CDPA was first implemented in 2011 at the University of British Columbia (UBC) and the University of Edinburgh [10]. It was initially developed for first-year physics major students in laboratory courses or in science-intensive programs that also required an experimental physics laboratory component. Graduate students and professors also participated in the study in order to ensure the validity of the test in measuring novice vs. expert performance and to further investigate its reliability when used by students at different levels of instruction. The CDPA test consists of 10 multiple choice questions, which vary in terms of difficulty, each with four possible answers. The questions cover a variety of learning outcomes such as identifying the adequacy of a fit to a given dataset, handling instrumental uncertainty, and inferring measurement units from the data. The test’s characteristics, such as its difficulty and its capacity to discriminate amongst the level of expertise of test-takers, are proposed and quantified in the original paper. Furthermore, the test is administered both as a pre-course and post-course assessment in order to reliably measure shifts in reasoning that can be linked to the learning outcomes of each course.
@@
The results in [10] showed 1) that there are positive shifts in pre- to post- comparisons for the first-year courses, 2) that the test is sensitive to the level of expertise of the test-taker, and thus is able to distinguish between undergraduate students, graduate students, and professors, 3) that the test has wide applicability, a conclusion made based on the comparable results from UBC and the University of Edinburgh. All these findings are supported by statistical tests which confirm the reliability of the CDPA. For example, first-year UBC students scored an average of $3.91\pm 0.12$ (standard error of the mean, SEM) in the post-test, compared to an average of $5.31\pm 0.42$ (SEM) for the graduate students, and $8.00\pm 0.35$ (SEM) for professors. A similar, more marginal progression of knowledge was also shown across the undergraduate program by comparing the post-test of first-year students with that of fourth-year students who averaged $4.73\pm 0.26$ (SEM) in the assessment.
@@
Since its publication, the CDPA has stimulated further research within the PER community with different applications of the CDPA results. For example, Day et al. [21] investigated the influence of gender in the assessment results, finding no evidence of gender differences in conceptual understanding of uncertainties. In addition, Holmes et al. [22] used the CDPA to investigate the role of guidance in productive failure during the learning process. Furthermore, Kontro [23] explored the use of the CDPA for intermediate-level undergraduate physics courses and proposed the possibility of a relationship between CDPA final scores and E-CLASS results, though no significant correlation is observed.
@@ III.1 Context
The research conducted for this study took place between 2019 to 2023. This period included the global pandemic in which instructors were forced to adopt new pedagogical strategies for the online environment, while maintaining course content and learning outcomes. When the pandemic hit in 2020, courses moved completely online. In 2021, a hybrid format was adopted, in which lectures were held online and lab sessions were mostly delivered in person. During 2022, lectures and labs transitioned back to a fully in-person format. In keeping with the CDPA requirements, no adaptations were made to the original CDPA test and the entire ten-question test was provided to all student respondents. Furthermore, as requested by the authors, we did not reproduce or include direct citations of any of the ten questions in order to preserve the use of this instrument for future research. To incentivize participation, a bonus of $1\%$ was offered to all students in each class if at least $80\%$ of students completed the CDPA test.
Finally, to avoid statistically significant shifts due to students retaking the CDPA test from one year to the next, we asked student respondents to identify whether they had taken the CDPA test in a previous semester. The test was conducted online through MS Forms and student respondents were instructed to use up to 30 minutes to complete it. The pre-tests were taken over the first three to four weeks of term, whereas post-tests were taken during the last three to four weeks of term. Tests without pre and post distinction were taken at the end of the term. We collected data in 8 courses: PHYS 101- Introduction to Mechanics, PHYS 131- Mechanics and Waves, PHYS 257- Experimental Methods I, PHYS 258- Experimental Methods II, PHYS 359- Honors Laboratory in Modern Physics I, PHYS 439- Majors Laboratory in Modern Physics, PHYS 469- Honors Laboratory in Modern Physics 2, and PHYS 434- Optics. All results presented here are summarized in Appendix appendix.
@@ III.2 McGill Physics curriculum
The understanding of uncertainty was identified by professors as a core concept in the undergraduate Physics laboratory curriculum. This is reflected in the increasing weight given to assessments that measure this concept as students move from introductory to advanced laboratory courses.
At McGill University, there are different streams for students taking lab courses: first-year students in the general science program (or those pursuing a minor in physics) are required to take an introductory undergraduate laboratory course, PHYS 101- Introduction to Mechanics; physics majors and honors students take a higher-level course PHYS 131- Mechanics and Waves. In both these courses, students explore basic mechanics concepts through simple confirmation experiments where the concept of uncertainty is addressed but does not constitute a major learning outcome. Students who proceed to a physics-specific degree, i.e., majors and honors students, enroll during their second year in PHYS 257- Experimental Methods I and PHYS 258- Experimental Methods II, where they are formally introduced to uncertainties in experiments. In these two more advanced courses, students are required to report uncertainties and discuss the validity of their experimental results and findings based on their interpretation of these uncertainties.
By the conclusion of PHYS 257 and PHYS 258, students are expected to possess the skills needed to interpret, address, and visually represent measurements. Students who choose to focus on experimental physics take the 300- and 400-level courses, which are not part of the mandatory curriculum, and can therefore represent expertise-like level of understanding. The notation difference between 300 and 400 is simply a means to indicate which students are part of the honors program as opposed to the majors program, otherwise the curricular content covered in both PHYS 359 and PHYS 439 is identical. We combined the results for advanced courses (PHYS 434, PHYS 439 and PHYS 469) to form the overall 400 level, because the sample size of each class was too small and the students’ level of expertise was similar.
@@ III.3 Educational background
In the context of our university, situated in Québec, a notable portion of incoming freshman students exhibit diverse educational backgrounds. This diversity is attributed to the fact that some students have completed Collège d’Enseignement Général et Professionnel (CEGEP) before enrolling in university. Students who complete CEGEP and pursue a physics degree at univeristy skip the first year of the McGill curriculum, taking PHYS 257 and PHYS 258 as their initial university-level laboratory courses. During data collection and analysis, we could not differentiate whether students had completed CEGEP or not.
@@ III.4.1 Origin of the PER group
The PER group leading this study represented a partnership between the Department of Physics and the Office of Science Education (OSE). Following an in-depth curriculum review process, the chair of the department nominated five tenure track professors teaching at different levels in the experimental physics program to work with OSE staff (including the director, educational developers and a science education expert), the physics liaison librarian, as well as PhD students selected by the department. The goal was to measure the effectiveness of lab courses from a program perspective. Although membership evolved over the five-year period, the group met on a regular basis and operated similarly to a community of practice [cuddy2002cultivating] with distributed leadership and regular cycles of data collection and reflection.
@@ III.4.2 PER professors feedback sessions
Following the most recent round of data collection, feedback sessions were held with the PER professors. The purpose of these sessions was to better understand the level of each department’s experimental courses and to map the specific skills taught in each course as they related to the content of the CDPA questions. These sessions were used to situate each course in the McGill Physics undergraduate program based on their content, and to better understand how students at various points in their degrees would be expected to respond to the CDPA. This analysis will be discussed further below.
@@ IV.1 Comparison to UBC
To validate the CDPA tool, in 2019 we decided to compare McGill’s CDPA results to those at University of British Columbia (UBC), where the tool was first developed. In Figure 1, we present the results at the two universities. The results demonstrate a similar difference between novice and expert conceptual understanding of uncertainties at both universities. We notice a maximum score for students in the fourth and final year of their undergraduate education, without significant change towards the graduate student level. As expected, professors from both universities score higher than all previous cohorts, but not a perfect score. These CDPA results allow us to draw an important conclusion: even at the higher end of the novice to expert scale, the average score is not a complete $10$, a feature consistent with Day et. al. [10].
@@ IV.2 Longitudinal analysis of the McGill data
The total CDPA mean score for the courses investigated between 2019-2023 is presented on a scale from 0 to 10 in Figure 2a. From left to right, the order of level of expertise increases by course numbers. We notice statistically significant jumps in the CDPA scores when comparing the scores obtained during the very first mechanics course that contains a lab component, PHYS 101, to a more advanced laboratory, PHYS 257. This can be further explained by the fact that in PHYS 257 students are more equipped to deal with uncertainties and apply methods to analyze them experimentally.
@@
In Figure 2b, we compare how the CDPA scores evolve for the same course as a function of time. Interestingly, for PHYS 101 and PHYS 131, the overall CDPA score has a combined average that does not change as a function of year, except in the rare case of PHYS 101 in 2022 where the score was abnormally low. It is important to note that no major changes were implemented in the course that could have explained this difference in score. In PHYS 257 and PHYS 258, the results show fluctuations that can be consistent with the learning curve faced at that level for each of the different cohorts. The structure, content and evaluation methods used in these two courses did not change in any of the years under study. However, different professors taught the course each year using the same syllabus. In the 2020 Winter session of PHYS 258, due to the pandemic, students only completed 60 % of the course – the laboratory component. This may partially explain the decrease in CDPA scores that year.
For higher level courses over the first three years, the overall score is also consistent with a jump observed in 2022. In that year, the number of students taking part in the class was less than in previous years but students still obtained significantly higher scores than previous cohorts. These findings are likely due to classroom dynamics that are notably different from the 200 level courses, which we will further argue plays an important role in students’ understanding of uncertainty.
@@
The data presented in Figure 2 shows that for the two lower-level courses, PHYS 101 and PHYS 131, the overall CDPA score can be understood as being comparable to random guessing. This can be explained by the fact that, in those courses, students are introduced to the concept of uncertainties for the first time and are only capable of applying simple mechanics problems and experiments without a deeper conceptual understanding. The CDPA test evaluates student understanding in topics that involve uncertainty propagation, graph analysis with fits and fitting uncertainties, logarithmic fits, and evaluation of $\chi^{2}$. In the original CDPA paper [10], the authors discuss that the vast majority of students from introductory-level laboratory courses are not yet ready to answer all the survey questions in the first year of their undergraduate degree. These results are in agreement with our observations and analysis of the understanding of uncertainty for students enrolled in PHYS 101 and PHYS 131.
In the following discussion, we hone in on the results obtained from the PHYS 257 and PHYS 258 courses because it is during this full academic year that students learn how to perform curve fitting, start considering uncertainties to assess the outcome of an experiment, and produce and interpret graphs based on measuring uncertainties. Comparisons at the higher level could also garner interesting results, however the population size of student respondents was not sufficiently large for each year of the study. Additionally, having pairwise comparisons between two courses within the same academic year allows for a better comparison in terms of the understanding of uncertainties. In the case of the McGill Physics curriculum, the two courses are complementary to one another and students’ abilities to understand concepts related to uncertainty is not developed until the end of the full academic year, as shown in Table 1. At the end of PHYS 258, students are asked to conduct a small research project in which they need to answer a research question of their choice. Students experimentally validate their measurements and study how the uncertainties should be considered in the analysis, as well as how to best mitigate errors.
@@
IV.3 Impact of the pandemic
The global COVID-19 pandemic of 2020 introduced a significant external factor that we were able to study. Before, during, and after the pandemic the content of the courses included in this study did not change. However, the delivery formats changed due to regulatory health guidelines. Once smaller courses returned to an in-person teaching environment, many of the large-scale introductory courses, including PHYS 101, PHYS 131, PHYS 257, and PHYS 258 continued to be taught in either a hybrid or a fully remote format.
Surprisingly, we notice no significant difference in CDPA scores between these pre- and post-pandemic periods. As was pointed out by one of the course professors, this lack of any statistically significant fluctuation in CDPA scores is, in and of itself, noteworthy. The consistency of these scores, as shown in Figures 2 and Figure 3, suggests that the pedagogical benefit of laboratory courses may have more to do with students gaining first-hand practical knowledge through manipulating laboratory instruments than with their conceptual understanding of measurement principles.
Furthermore, if exposure to experimental instruments is only part of the intended pedagogical value of laboratory courses, then how can these in-person labs be improved such that they enhance students’ understanding of concepts related to uncertainty to a degree that is distinguishable using a measurement instrument such as the CDPA? The consistency of these pandemic-era CDPA scores also raises the question whether the CDPA instrument can distinguish between the impact of confirmation versus inquiry-guided instructional approaches. From the findings of this research, we deduce that in all undergraduate courses included in this study could benefit from more dedicated teaching of graphing and analysis of uncertainties, as we will discuss in the following sections.